# Development of a conditional diffusion model to predict process parameters and microstructures of dendrite crystals of matrix resin based on mechanical properties

### v1

@@ 3.3 Results for conditions not included in the dataset
We confirm the results for the conditions not included in the dataset. The results are shown in Figs. 1520. Fig. 15 shows the generated images of microstructures and Fig. 16 shows the generated images that indicate the processing temperature. Fig. 17 shows the relationship between the proposed temperature and the elasticity matrix $\boldsymbol{D}$ obtained by homogenization analysis using XFEM for the generated microstructure. Fig. 18 is a scatter plot with $D_{1111}$ on the horizontal axis and $D_{2222}$ on the vertical axis, plotted in different colors for different crystallization temperatures. Figs. 17 and 18 show that the relationship between crystallization temperature and the elasticity matrix $\boldsymbol{D}$ is generally similar to that of the training and test data, even when the data are generated using the conditions not included in the dataset. Fig. 19 shows examples of the generated results. The model developed outputs images of microstructures and IPPTs. Fig. 19 shows (a) the generated microstructures, (b) IPPTs, (c) the conditions and (d) the elasticity matrix $\boldsymbol{D}$ values obtained by homogenization analysis using XFEM for the generated microstructure. As shown in Fig. 19(a), the conditional diffusion model can represent the detailed microstructure of dendrite crystals of matrix resins. By homogenization analysis with XFEM to obtain the elasticity matrix $\boldsymbol{D}$ for the predicted microstructure, we find that elasticity matrix $\boldsymbol{D}$ is close to the condition given for generation. Fig. 20 shows the elasticity matrix $\boldsymbol{D}$ for each component. The horizontal axis represents the elasticity matrix $\boldsymbol{D}$ obtained by homogenization analysis using XFEM for the generated microstructure, and the vertical axis is the condition given for generation. The correlation coefficients and each plot indicate that the values of elasticity matrix $\boldsymbol{D}$ obtained by homogenization analysis using XFEM agree with the condition given for generation with good accuracy. The correlation coefficients imply a strong positive correlation between input condition and calculated result by XFEM, suggesting appropriate microstructure for the given condition.
@@ 3.4 Demonstration of the proposed conditional diffusion model
The conditional diffusion model developed in this study can propose the optimal processing temperature and predict the microstructure for given Young’s modulus and Poisson’s ratio. The advantage of this model is that it outputs not only the processing temperature but also the microstructure as images. Fig. 21 shows example of the model demonstration. Fig. 21 shows (a)(i) the specified values of Young’s modulus, (a)(ii) the specified values of Poisson’s ratio, (b)(i) the proposed temperatures and (b)(ii) the predicted microstructures of matrix resins. We train the model with the conditions obtained by homogenization analysis using XFEM on $320{\ \mathrm{pixels}}\times 320{\ \mathrm{pixels}}$ before image compression. Table 5 shows the number and size of data used for training. The relationship between Young’s modulus and Poisson’s ratio and the calculation of the elasticity matrix $\boldsymbol{D}$ is described in A. As shown in Fig. 21, 200 C is proposed for a large Young’s modulus of $2,761\,\text{MPa}$, whereas 160 C is suggested for a small Young’s modulus of $2,210\,\text{MPa}$. Thus, the results show that the temperature tends to increase with Young’s modulus.
@@ 4 Conclusions
In this study, we developed a conditional diffusion model that can propose the optimal processing temperature and predict the microstructure for the desired elastic constants. As an example, we analyzed PPS, a thermoplastic resin that is the base material of CFRTP, under the assumption of composite material applications. As a result of using the developed model, the following three were confirmed:
  1. 1.
@@
In conclusion, the conditional diffusion model developed in this study can propose the optimal processing temperature and predict the microstructure of matrix resins that satisfies the desired mechanical properties. This study enables us to realize an inverse design that proposes the process parameters related to how the material should be made so as to satisfy the mechanical properties. The model can be extended to composites by including fibers in the analysis. By replacing the training data of microstructure, process temperature and elasticity matrix, this model can be applied to other materials such as CFRTS containing polymer blends with phase-separated structures.
Acknowledgments
@@
Appendix A Homogenization analysis and elasticity matrix

### v2

@@ 3.3 Results for conditions not included in the dataset
We confirm the results for the conditions not included in the dataset. The results are shown in Figs. 1621. Fig. 16 shows the generated images of microstructures and Fig. 17 shows the generated images that indicate the processing temperature. Fig. 18 shows the relationship between the proposed temperature and the elasticity matrix $\boldsymbol{D}$ obtained by homogenization analysis using XFEM for the generated microstructure. Fig. 19 is a scatter plot with $D_{1111}$ on the horizontal axis and $D_{2222}$ on the vertical axis, plotted in different colors for different crystallization temperatures. Figs. 18 and 19 show that the relationship between crystallization temperature and the elasticity matrix $\boldsymbol{D}$ is generally similar to that of the training and test data, even when the data are generated using the conditions not included in the dataset. Fig. 20 shows examples of the generated results. The model developed outputs images of microstructures and IPPTs. Fig. 20 shows (a) the generated microstructures, (b) IPPTs, (c) the conditions and (d) the elasticity matrix $\boldsymbol{D}$ values obtained by homogenization analysis using XFEM for the generated microstructure. As shown in Fig. 20(a), the conditional diffusion model can represent the detailed microstructure of dendrite crystals of thermoplastic resins. By homogenization analysis with XFEM to obtain the elasticity matrix $\boldsymbol{D}$ for the predicted microstructure, we find that elasticity matrix $\boldsymbol{D}$ is close to the condition given for generation. Fig. 21 shows the elasticity matrix $\boldsymbol{D}$ for each component. The horizontal axis represents the elasticity matrix $\boldsymbol{D}$ obtained by homogenization analysis using XFEM for the generated microstructure, and the vertical axis is the condition given for generation. The correlation coefficients and each plot indicate that the values of elasticity matrix $\boldsymbol{D}$ obtained by homogenization analysis using XFEM agree with the condition given for generation with good accuracy. The correlation coefficients imply a strong positive correlation between input condition and calculated result by XFEM, suggesting appropriate microstructure for the given condition.
@@ 3.4 Demonstration of the proposed conditional diffusion model
The conditional diffusion model developed in this study can propose the optimal processing temperature and predict the microstructure for given Young’s modulus and Poisson’s ratio. The advantage of this model is that it outputs not only the processing temperature but also the microstructure as images. Fig. 22 shows example of the model demonstration. Fig. 22 shows (a)(i) the specified values of Young’s modulus, (a)(ii) the specified values of Poisson’s ratio, (b)(i) the proposed temperatures and (b)(ii) the predicted microstructures of thermoplastic resins. We train the model with the conditions obtained by homogenization analysis using XFEM on $320{\ \mathrm{pixels}}\times 320{\ \mathrm{pixels}}$ before image compression. Table 5 shows the number and size of data used for training. The relationship between Young’s modulus and Poisson’s ratio and the calculation of the elasticity matrix $\boldsymbol{D}$ is described in Homogenization analysis and elasticity matrix. As shown in Fig. 22, 200 C is proposed for a large Young’s modulus of $2,761\,\text{MPa}$, whereas 160 C is suggested for a small Young’s modulus of $2,210\,\text{MPa}$. Thus, the results show that the temperature tends to increase with Young’s modulus.
@@ 4 Conclusions
In this study, we developed a conditional diffusion model that can propose the optimal processing temperature and predict the microstructure for the desired elastic constants of thermoplastic resins. As a result of using the developed model, the following three were confirmed:
@@
In conclusion, the conditional diffusion model developed in this study can propose the optimal processing temperature and predict the microstructure of thermoplastic resins that satisfies the desired mechanical properties. This study enables us to realize an inverse design that proposes the process parameters related to how the material should be made so as to satisfy the mechanical properties.
This model can be applied to other materials, process parameters and mechanical properties by replacing the data used for training, such as the microstructure, processing temperature and elastic constants.
Acknowledgements
@@
Conflict of interest
The authors declare that they have no conflicts of interest to disclose.
Contributions
A.I. conducted Data curation, Formal analysis, Investigation, Methodology, Software, Visualization, Validation, Writing - Original Draft, Writing - Review and Editing. R.H. conducted Conceptualization, Funding acquisition, Investigation, Methodology. T.Y. conducted Conceptualization, Funding acquisition, Methodology. K.E. conducted Conceptualization, Investigation, Methodology, Software, Validation. Y.K. conducted Investigation, Methodology, Software, Validation, Writing - Review and Editing. M.S. conducted Investigation, Methodology, Software, Validation. M.M. conducted Conceptualization, Funding acquisition, Project administration, Resources, Writing - Review and Editing, Supervision.
Appendix
Homogenization analysis and elasticity matrix